# Community-Centric Graph Unlearning

### v1

@@ Abstract
Graph unlearning technology has become increasingly important since the advent of the ‘right to be forgotten’ and the growing concerns about the privacy and security of artificial intelligence. Graph unlearning aims to quickly eliminate the effects of specific data on graph neural networks (GNNs). However, most existing deterministic graph unlearning frameworks follow a balanced partition-submodel training-aggregation paradigm, resulting in a lack of structural information between subgraph neighborhoods and redundant unlearning parameter calculations. To address this issue, we propose a novel Graph Structure Mapping Unlearning paradigm (GSMU) and a novel method based on it named Community-centric Graph Eraser (CGE). CGE maps community subgraphs to nodes, thereby enabling the reconstruction of a node-level unlearning operation within a reduced mapped graph. CGE makes the exponential reduction of both the amount of training data and the number of unlearning parameters. Extensive experiments conducted on five real-world datasets and three widely used GNN backbones have verified the high performance and efficiency of our CGE method, highlighting its potential in the field of graph unlearning.
@@ Introduction
As Artificial Intelligence (AI) permeates various sectors, protecting individual privacy and data security has become an increasingly pressing concern. Legal frameworks, such as the “right to be forgotten” (Pardau 2018; Mantelero 2013; Cofone 2020), alongside the rise of techniques like membership inference attacks (He et al. 2021a) have underscored the need for model providers to swiftly remove specific data and mitigate its impacts. In parallel, Graph Neural Networks (GNNs) (Li et al. 2022; Xu et al. 2018) have gained prominence as a powerful tool, renowned for their ability to model complex data relationships. However, their widespread application in areas involving highly sensitive data, such as social networks(Qiu et al. 2018) and recommendation systems (Fan et al. 2019), has intensified the demand for effective privacy-preserving techniques. Among these, Graph Unlearning (Chen et al. 2022b) has gained particular prominence.
@@
where $\mathcal{G}^{\prime}=(\mathcal{V}\setminus\mathcal{V}_{u},\mathcal{E}\setminus E_{u})$ and $\theta^{\prime}$ denotes the updated model parameters after unlearning. The advanced objectives of graph unlearning should ensure portability, comparable model utility to training from scratch (Chen et al. 2022b), high graph structure utilization, and minimal unlearning window periods. Graph unlearning can be broadly divided into node-level and edge-level approaches. In our GSMU paradigm, edge unlearning is considered a subproblem of node unlearning. Consequently, our work focuses on the node classification task, treating node unlearning as the fundamental unit.
@@
$\widetilde{\mathcal{G}}^{\prime}=\operatorname{Update_{\mathcal{\widetilde{G}}}}(\widetilde{\mathcal{V}}^{\prime},\widetilde{\mathcal{E}}^{\prime},\widetilde{\mathcal{X}}^{\prime},\widetilde{\mathcal{Y}}^{\prime}),$ (14)
@@
$f_{\theta^{\prime}}(\widetilde{\mathcal{G}}^{\prime})\approx f_{\theta}(\mathcal{\widetilde{G}})\quad\text{subject to}\quad\mathcal{V}_{u}\notin\widetilde{\mathcal{G}}^{\prime},$ (15)
@@
Table 4: The comparison of two-stage time consumption of different graph unlearning frameworks on four datasets(s).
@@ Results.
The experimental results presented in Table 3 reveal the following findings: (1) CGE outperforms all established baselines in terms of model utility. For instance, on Reddit, CGE shows an average increase of 11.56% compared to GUIDE and an even more significant average increase of 33.24% compared to GraphEraser. This superiority is attributed to two factors: (a) CGE employs a graph-optimized community detection method, resulting in a more accurate community distribution than the balanced partitioning used in the BP-SM-TA paradigm. (b) CGE accounts for the inherent relationships between subgraphs, reducing information loss, a consideration overlooked by the BP-SM-TA paradigm. (2) CGE demonstrates comparable model utility to Scratch, which serves as the foundational baseline due to its broad and outstanding performance achieved through the entire retraining of the dataset. Notably, CGE has even outperformed Scratch in certain cases, primarily due to its pre-integration of similar data within the same community in the original graph, effectively reducing noise. Consequently, the newly mapped graph in CGE contains more robust and accurate information than the original graph.
@@
To determine the effectiveness of CGE in unlearning, we utilize a node-level GNN member inference attack methodology (MIA) (He et al. 2021b). This approach involves calculating the discrepancy between the original model and the model after CGE unlearning, to determine whether a node has been successfully removed from the model. After randomly removing 0.5% of all nodes according to the experimental parameters, the MIA attack was executed. We consider two scenarios: $\mathcal{A}$ and $\mathcal{A_{C}}$, representing the MIA attack results on the original GNN backbones and CGE, respectively. As shown in Table 6, the AUC of MIA on CGE is significantly lower than that of the original GNN backbones, approaching 0.5, which is analogous to random guessing. This indicates that CGE does not produce additional information leakage and that its unlearning of nodes and associated effects is effective.
@@ Conclusions
While CGE shows great potential, there are areas for improvement. Traditional community detection techniques may prove inadequate when applied to large-scale graphs. Our ongoing research into deep learning-based community detection methods aims to balance the privacy implications of deep learning with its potential for enhancing data insights. This represents a critical area of focus for our future work.

### v2

@@ Abstract
Graph unlearning technology has become increasingly important since the advent of the ‘right to be forgotten’ and the growing concerns about the privacy and security of artificial intelligence. Graph unlearning aims to quickly eliminate the effects of specific data on graph neural networks (GNNs). However, most existing deterministic graph unlearning frameworks follow a balanced partition-submodel training-aggregation paradigm, resulting in a lack of structural information between subgraph neighborhoods and redundant unlearning parameter calculations. To address this issue, we propose a novel Graph Structure Mapping Unlearning paradigm (GSMU) and a novel method based on it named Community-centric Graph Eraser (CGE). CGE maps community subgraphs to nodes, thereby enabling the reconstruction of a node-level unlearning operation within a reduced mapped graph. CGE makes the exponential reduction of both the amount of training data and the number of unlearning parameters. Extensive experiments conducted on five real-world datasets and three widely used GNN backbones have verified the high performance and efficiency of our CGE method, highlighting its potential in the field of graph unlearning. https://github.com/liiiyi/CCGU
@@ Introduction
As Artificial Intelligence (AI) permeates various sectors, protecting individual privacy and data security has become an increasingly pressing concern (Wei et al. 2024). Legal frameworks, such as the “right to be forgotten” (Pardau 2018; Mantelero 2013; Cofone 2020), alongside the rise of techniques like membership inference attacks (He et al. 2021) have underscored the need for model providers to swiftly remove specific data and mitigate its impacts. In parallel, Graph Neural Networks (GNNs) (Li et al. 2022; Xu et al. 2018; Li et al. 2024a, b) have gained prominence as a powerful tool, renowned for their ability to model complex data relationships. However, their widespread application in areas involving highly sensitive data (Zhang et al. 2025), such as social networks(Qiu et al. 2018) and recommendation systems (Fan et al. 2019), has intensified the demand for effective privacy-preserving techniques. Among these, Graph Unlearning (Chen et al. 2022b) has gained particular prominence.
@@
where $\mathcal{G}^{\prime}=(\mathcal{V}\setminus\mathcal{V}_{u},\mathcal{E}\setminus E_{u})$ and $\theta^{\prime}$ denotes the updated model parameters after unlearning. The advanced objectives of graph unlearning should ensure portability, comparable model utility to training from Scratch (Chen et al. 2022b), the maintenance of model performance before and after unlearning, high graph structure utilization, and minimal unlearning window periods. Graph unlearning can be broadly divided into node-level and edge-level approaches. In our GSMU paradigm, edge unlearning is considered a subproblem of node unlearning. Consequently, our work focuses on the node classification task, treating node unlearning as the fundamental unit.
@@
$\widetilde{\mathcal{G}}^{\prime}=\operatorname{Update_{\mathcal{\widetilde{G}}}}(\widetilde{\mathcal{V}}^{\prime},\widetilde{\mathcal{E}}^{\prime},\widetilde{\mathcal{X}}^{\prime},\widetilde{\mathcal{Y}}^{\prime}).$ (14)
@@
$f_{\theta^{\prime}}(\widetilde{\mathcal{G}}^{\prime})\approx f_{\theta}(\mathcal{\widetilde{G}})\quad\text{subject to}\quad\mathcal{V}_{u}\notin\widetilde{\mathcal{G}}^{\prime}.$ (15)
@@
Table 4: The comparison of two-stage time consumption of different graph unlearning frameworks on four datasets (second).
@@ Results.
The experimental results presented in Table 3 reveal the following findings: (1) CGE outperforms all established baselines in terms of model utility. For instance, on Reddit, CGE shows an average increase of 11.56% compared to GUIDE and an even more significant average increase of 33.24% compared to GraphEraser. This superiority is attributed to two factors: (a) CGE employs a graph-optimized community detection method, resulting in a more accurate community distribution than the balanced partitioning used in the BP-SM-TA paradigm. (b) CGE accounts for the inherent relationships between subgraphs, reducing information loss, a consideration overlooked by the BP-SM-TA paradigm. (2) CGE demonstrates comparable model utility to Scratch, which serves as the foundational baseline due to its broad and outstanding performance achieved through the entire retraining of the dataset. Notably, CGE has even outperformed Scratch in certain cases, primarily due to its pre-integration of similar data within the same community in the original graph, effectively reducing noise. This results in a newly mapped graph with more robust and accurate information. CGE achieves average performance closer to Scratch, fulfilling the objectives of unlearning.
@@
To determine the effectiveness of CGE in unlearning, we utilize a node-level GNN member inference attack methodology (MIA) (He et al. 2021). This approach involves calculating the discrepancy between the original model and the model after CGE unlearning, to determine whether a node has been successfully removed from the model. After randomly removing 0.5% of all nodes according to the experimental parameters, the MIA attack was executed. We consider two scenarios: $\mathcal{A}$ and $\mathcal{A_{C}}$, representing the MIA attack results on the original GNN backbones and CGE, respectively. As shown in Table 6, the AUC of MIA on CGE is significantly lower than that of the original GNN backbones, approaching 0.5, which is analogous to random guessing. This indicates that CGE does not produce additional information leakage and that its unlearning of nodes and associated effects is effective.
@@ Conclusions
CGE shows strong potential but faces challenges with traditional community detection techniques on large-scale graphs. Future work will leverage deep learning-based methods to enhance data insights while addressing privacy concerns, aiming for a balanced and effective approach.
Acknowledgments
This work was partially supported by the Project of Guangxi Science and Technology (GuiKeAB23026040), Research Fund of Guangxi Key Lab of Multi-source Information Mining & Security (MIMS24-13), Research Fund of Guangxi Key Lab of Multi-source Information Mining & Security (24-A-01-02), and Innovation Project of Guangxi Graduate Education (XYCSR2024102).
@@ Acknowledgments
Cora(M) Cora(O) Citeseer(M) Citeseer(O) CS(M) CS(O) Reddit(M) Reddit(O)
Entropy 1.8197 1.8311 1.7434 1.7533 2.3961 2.4087 3.3967 3.4013
IR 0.2024 0.2200 0.3068 0.3766 0.0291 0.0285 0.0141 0.0116
Gini 0.2706 0.2587 0.1400 0.1338 0.4328 0.4225 0.4314 0.4222
Table 8: Class distribution and imbalance measures for mapped and original graphs. (M) denotes the mapped graph generated by CGE and (O) denotes the original graph.
@@
Appendix E E. Evaluation of Class Imbalance in CGE
In the framework design, it was considered that some nodes within a single community may represent noise at the feature or label level relative to the entire collective. To mitigate the complete dominance of the majority class, an intuitive approach was proposed based on majority voting and Euclidean distance-based local proximity. This allows label selection to better align with the community’s local structure, thereby minimizing the negative impact on supervised learning.
To validate the approach, we conducted experiments involving the class distribution of nodes in the generated mapped graph and the original graph. The class imbalance was quantified using three mathematical measures: class distribution vector entropy, class imbalance ratio (IR), and the Gini coefficient. The following outlines the methods used for calculating these measures:
  • Class Distribution Vector Entropy: The entropy of the class distribution vector was calculated using the formula:
    $H=-\sum_{i=1}^{n}p_{i}\log(p_{i}),$ (28)
    where $p_{i}$ is the proportion of nodes belonging to class $i$, and $n$ is the number of distinct classes. This measure quantifies the uncertainty or disorder within the class distribution.
  • Class Imbalance Ratio (IR): The class imbalance ratio is defined as the ratio between the size of the largest class and the size of the smallest class:
    $IR=\frac{\text{Size of Largest Class}}{\text{Size of Smallest Class}}.$ (29)
    A higher ratio indicates greater imbalance in the class distribution.
  • Gini Coefficient: The Gini coefficient is calculated as:
    $G=1-\sum_{i=1}^{n}p_{i}^{2},$ (30)
    where $p_{i}$ is the proportion of nodes in class $i$. This measure provides a summary of the inequality of the class distribution, with higher values indicating greater imbalance.
The results of the experiment are summarized in Table 8, where (M) denotes the mapped graph generated by CGE and (O) denotes the original graph.
The results indicate that the mapped graph generated by CGE did not introduce additional data bias when compared to the original graph.